OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (-1,1) and (3,-3)

OpenStudy (anonymous):

Slope-intercept: Y = MX+B... First, solve for M. M= (y2-y1) / (x2-x1) Set a point as each ONE and TWO... In this case, I will use pt1 is (-1,1) and pt2 is (3,-3). Plugging in, I get (-3 - 1) / (3- -1). Which becomes -4 / 4. M= -1. Now, back in our original equation, we plug in -1 for M. Y=-1(X) +B. We use a point (either, it doesn't matter which) to determine B... In this case, I will use (-1, 1)... Thus, 1=(-1)(-1) +B. 1=1+B... B=0. So, our final answer is Y=-1X